OpenStudy (anonymous):

does any 1 know how twe do this 24+z=43

Parth (parthkohli):

Can you subtract both sides by 24?

OpenStudy (anonymous):

subtract 24 on both sides. then divide

Parth (parthkohli):

\(\begin{array} {}24 + z = 43 \\ -24 \quad -24 \\ \hline z = 43 - 24 \end{array}\)

OpenStudy (anonymous):

24+x-24=43-24 x=19

OpenStudy (anonymous):

z=43-24

OpenStudy (anonymous):

thx so much u guys helped alot

Parth (parthkohli):

Can you tell me what 43 - 24 is?

OpenStudy (anonymous):

i think it is 19

Parth (parthkohli):

Yep!

OpenStudy (anonymous):

kk thx for the help so tht is the answer rite

OpenStudy (amistre64):

the answer is the result of the process ....